Jacks sales sells photocopiers and provides parts and services to its customers. The company has 30 employees and provide employees with a benefit plan. You are the controller and are looking at the liability section of the balance sheet and have identified several areas for consideration.

  1. The company remits HST quarterly. The company had sales of $2,400,000 (before HST) and made purchases of $1,300,000 (before HST) during the quarter. The HST rate is 15%. A customer retuned a sale in the amount of $3,500 plus HST.
    1. Record the journal entries for the sales, purchases and remittance of HST for the quarter. Assume the company uses a periodic inventory system.

  1. The payroll for December was as follows: Total payroll $170,000 all of which is considered pensionable and insurable, income taxes in the amount of $44,800 were withheld as well as $1,600 in union dues. CPP rate for employees and employers is 4.95%. EI rate for employees is 1.66% and employers must pay 1.4 times the employees rate.

  1. Prepare the December journal entries including the remittance.
Answer #1

For Part A Journal Entries are-

A. For Sales-

Vendor Account/Cash/Bank Debit. $2,396,500

Sales Return Account Debit $3500

To Sales Account $ 2,400,000

B. For Purchase-

Purchase Account Debit $1100000

To Cash/Bank/Vendor Account $1,100,000

C. For HST Liability-

Profit and loss Account Debit $194,475

To Provision for HST liability Account $194,475

At the time of HST Payment-

Provision for HST liability Account Debit $194,475

To Cash/ Bank Account $194,475

HST Liability= (2,400,000-3,500-11,00,000)*15%

Part B Journal Entries-

Salaries Account Debit $170,000

To Provision for Income tax on Salaries $44800

To Provision for Union Dues $1,600

To Provision for CPP(employees) $8,415

To Provision for CPP(employers) $8,415

To Provision for EI (employees) $2,822

To Provision for EI (employers) $3,951

To Cash/Bank Account $99,997

CPP= 170,000*4.95%=8,415

EI= For Employee 170,000*1.66%=2.822

For Employes 1,70,000*1,66%*1,4=3,950.8